What are the costs to the government of a subsidy if the demand and supply of Chinese coal are
Qd = 60 − 0.5P.
Qs = −2 + 2P.
1
Expert's answer
2020-11-12T10:35:25-0500

60-0.5p=-2+2p

2.5p=63

p=24.8

Q=60-0.5*24.8

Q=47.6